Using MATLAB

1. Assume Y is an exponential random variable with rate parameter λ=2. (1) Generate 1000 samples from this exponential distribution using inverse transform method (2) Compare the histogram of your samples with the true density of Y.
Answer #1

Assume \(Y\) is an exponential random variable with rate parameter \(\lambda = 2\). Recall that the probability density function is \(p(y) = 2e^{-2y}\), for \(y > 0\). First, we compute the CDF: \[F_Y(x) = P(Y\leq x) = \int_0^x 2e^{-2y} dy = 1 - e^{-2x}\]

Solving for the inverse CDF, we get that \[F_Y^{-1}(y) = -\frac{\ln(1-y)}{2}\]

Using our algorithm above, we first generate \(U \sim \text{Unif}(0,1)\), then set \(X = F_Y^{-1}(U) = -\frac{\ln(1-U)}{2}\). We do this in the R code below and compare the histogram of our samples with the true density of \(Y\).

# inverse transform sampling
num.samples <-  1000
U           <-  runif(num.samples)
X           <- -log(1-U)/2

# plot
hist(X, freq=F, xlab='X', main='Generating Exponential R.V.')
curve(dexp(x, rate=2) , 0, 3, lwd=2, xlab = "", ylab = "", add = T)

Generating Exponential R.V 0.0 0.5 1.0 1.5 2.0 2.5 3.0

Past versions of unnamed-chunk-2-1.png

Indeed, the plot indicates that our random variables are following the intended distribution